If daily demand is normally distributed with a mean of 15 and standard deviation of 5, and lead time is constant at 7 days, a 95% service level will require how much safety stock?
  • 10 units 
  • 13 units 
  • 17 units
  • 22 units
  • 8 units
Answer #1

Z = 1.645 (for 95% service level)

Lead time = 7 days

Standard deviation = 5 days

ROP = Z * standard deviation * sqrt (lead time)

= 1.645 * 5 * SQRT(7) = 21.76 or 22.

The correct option is 22 units.

Do give a positive rating if you think this answer helped